Required information

[The following information applies to the questions displayed below.]

On October 1, 2021, the Allegheny Corporation purchased equipment for $323,000. The estimated service life of the equipment is 10 years and the estimated residual value is $4,000. The equipment is expected to produce 550,000 units during its life.

Required:
Calculate depreciation for 2021 and 2022 using each of the following methods. Partial-year depreciation is calculated based on the number of months the asset is in service.

1. Straight line.
Answer #1

Calculate depreciation

2021 2022
Depreciation expense (323000-4000/10*3/12) = 7975 (323000-4000/10) = 31900